a. The moment of inertia of the reaction wheels has a value RI and the moment of inertia of the telescope about its symmetry axis, and about an axis perpendicular to its symmetry axis is represented by Iz and Ix , respectively. Find an expression for the final speed of the appropriate reaction wheel wRf to cause the telescope to change its direction at a rate of wt for a nominal (i.e. idling) rotational speed of the flywheels of wRi . Assume the initial rotational speed of the telescope is zero.

b. The moment of inertia of a solid cylindrical body (which is a passable approximation of a space telescope) is given by Iz=(1/2)mR^2 and Ix=(m/12)(R^2+H^2). Consider a space telescope of mass 1000 kg, radius 1.20 m and height 4.50 m and a moment of inertia of the reaction wheels of 3x10^-3 kgm^2 and an idling speed of the flywheels of 200 rad/s. Calculate the changes in the rotational speed required to cause a rotation of 0.100 degree per second for rotation
i. about the symmetry axis and
ii. about the axis perpendicular to the symmetry axis.
Answer #1

(b) Using a conservation of angular momentum, we have

L1 = L2

IT\omegaT = IW\omegaW

(i) about the symmetry axis -

IT = Iz = (1/2) m R2

[(0.5) (1000 kg) (1.2 m)2] \omegaT = (3 x 10-3 kg.m2) (200 rad/s)

(720 kg.m2) \omegaT = 600 kg.m2/s

\omegaT = 0.833 rad/s

The changes in the rotational speed required to cause a rotation of 0.1 degree per second for rotation which will be given as :

\Delta\omega = (0.8333 rad/s) - (0.0017 rad/s)

\Delta\omega = 0.831 rad/s

(ii) about the axis perpendicular to the symmetry axis -

IT = Ix = (1/12) m (R2 + H2)

(1/12) (1000 kg) [(1.2 m)2 + (4.5 m)2] \omegaT = (3 x 10-3 kg.m2) (200 rad/s)

(1807.5 kg.m2) \omegaT = 600 kg.m2/s

\omegaT = 0.3319 rad/s

The changes in the rotational speed required to cause a rotation of 0.1 degree per second for rotation which will be given as :

\Delta\omega = (0.3319 rad/s) - (0.0017 rad/s)

\Delta\omega = 0.33 rad/s